OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5).?

OpenStudy (diyadiya):

\[Slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (diyadiya):

\[\frac{-5-(-3)}{3-1}\]\[ \frac{-2}{2}\]\[=-1\]

OpenStudy (anonymous):

whats the slope of this gragh

OpenStudy (karatechopper):

-1

OpenStudy (karatechopper):

o tht one is.. u find out!

OpenStudy (karatechopper):

we gave u the formula!

OpenStudy (karatechopper):

now u plug the numbers in

OpenStudy (diyadiya):

The points are (-1.5,0) & (0,-3)

OpenStudy (karatechopper):

pick 2 points and use the formula

OpenStudy (anonymous):

ok you pick any two points?

OpenStudy (karatechopper):

yep

OpenStudy (diyadiya):

You cant take any two points You've to choose the points from the graph given Here as you can see points are (-1.5,0) & (0,-3)

OpenStudy (karatechopper):

thts what i meant didi!

OpenStudy (diyadiya):

Yes:)

OpenStudy (karatechopper):

:D

OpenStudy (diyadiya):

So now find the Slope using the Formula

OpenStudy (karatechopper):

Did my didi sweep twight?

OpenStudy (diyadiya):

\[Slope= \frac{-3-0}{0-(-1.5)}\]\[= \frac{-3}{1.5}=-2\]

OpenStudy (karatechopper):

lol

OpenStudy (anonymous):

so for this graph would the points be (7.5,0) (3,-3)?

OpenStudy (anonymous):

nd you do the same formula?

OpenStudy (diyadiya):

The points are (7.5,0) & (0,-5) (the line is passing through -5 not -3) Yes,Same formula for everything!

OpenStudy (anonymous):

okay i got -5 ----- -7.5 what do i do next?

OpenStudy (diyadiya):

\[=\frac{50}{75}\]\[= \frac{10}{15}\]\[=\frac{2}{3}\]

OpenStudy (diyadiya):

Just simplifiaction.

OpenStudy (anonymous):

ok i did what you said for this question Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). nd i got -5/-9 what did i do wrong ?